OpenStudy (anonymous):

Simplify sin^3(x)sec^2(x)-sin^2(x)

Nnesha (nnesha):

take out common factor

OpenStudy (anonymous):

there isn't one.

Nnesha (nnesha):

yes there is let's say sin = x and sec = y so x^3 y - x^2 what is common factor ?

OpenStudy (anonymous):

sin^2

Nnesha (nnesha):

yep right sin^2( ) what should be in the parentheses ?? :-)

OpenStudy (anonymous):

\[\sin^2x(sinxsec^2x-1)\]

OpenStudy (anonymous):

I need to simplify it more. How would I do that?

Nnesha (nnesha):

gimme a sec :3 i'm working on it

OpenStudy (anonymous):

ok. :) I've tried a lot of different ways myself and it never gets real simple.

OpenStudy (freckles):

doesn't really seem like you can simplify it

Nnesha (nnesha):

i guess try this identity 1+tan^2 =sec^2 otherwise that's it sin^2(sin(x)sec^2-1)

OpenStudy (freckles):

for some reason I read sin^2(x)sec^2(x)-sin^2(x) at first sin^2(x)(sec^2(x)-1) sec^2(x)tan^2(x)

OpenStudy (freckles):

but we had sin^3(x)sec^2(x)-sin^2(x)

Nnesha (nnesha):

ye

OpenStudy (freckles):

maybe we are just suppose to rewrite it (not really simplify) are there are choices @AEB047 ?

OpenStudy (anonymous):

no choices.

Nnesha (nnesha):

or you can write sin^3 as sin^2 times sin

OpenStudy (freckles):

earlier you said you needed to simplify more how did you know that like is this something you are trying to simplify to exactly?

Nnesha (nnesha):

\[ \sin^3(x)\sec^2(x)-\sin^2(x)\] write sec^2 = 1/cos^2 \[\sin(x) \times \sin^2(x) \times \frac{ 1 }{ \cos^2(x) } - \sin^2(x)\] sin^2/cos^2 = tan^2

OpenStudy (anonymous):

ok, so we can end up with: \[\frac{ \sin^3x }{ \cos^2x }-\sin^2x\] or \[\sin^2x(\frac{ sinx }{ \cos^2x }-1)\] or \[\tan^2xsinx-\sin^2x\] Which one is the most simplified, do you think?

OpenStudy (freckles):

it doesn't matter they are both ugly :p

Nnesha (nnesha):

i'll go with de last one :P i hate fraction >.<

OpenStudy (anonymous):

same. ok. thanks!

Nnesha (nnesha):

np :-)

OpenStudy (freckles):

@AEB047 can I ask where you got this question from: Simplify sin^3(x)sec^2(x)-sin^2(x)

OpenStudy (freckles):

Like did your teacher come up with it? Was that the original problem?

OpenStudy (freckles):

I'm just not convinced the "original problem" is actually the original problem Because to me no simpler form has been given to the "original" unless your teacher has clearly defined the word "simplify"